If you have a solution that's 16% alcohol and another that's 96% alcohol how much of the 96% alcohol do you have to add to the 16% to make the 16% raise to 18% alcohol?
Answer by stanbon(75887)

If you have 1000 ml of a solution that's 16% alcohol and another that's 96% alcohol how much of the 96% alcohol do you have to add to the 16% to make the 16% raise to 18% alcohol?
-------------------
Equation:
alch + alch = alch
-----
0.16*1000 +0.96*x = 0.18(1000+x)
-----
16*1000 + 96x = 18*1000 + 18x
78x = 2*1000
x = 25.64ml (amt. of 96% solution needed)
====================================
Cheers,
Stan H.